OpenStudy (sugarbloom2004):

I will be getting temporary zeroes in the subject pretty soon on the twenty-seventh of May, and I need someone such as a qualified tutor to help me with the quiz. I will medal anyone who gives me the right answer and walks me through the subject for the ten questions. The table shows the results of spinning a four-colored spinner fifty times. Find the experimental probability and express it as a decimal. P(not red) Red=20, Blue= 10, Green=9, Yellow=11. A.) 0.6 B.) 0.4 C.) 0.2 D.) 0.3. All spins amount to fifty, I divided the fifty by the colors that were not red.

satellite73 (satellite73):

right

OpenStudy (sugarbloom2004):

I chose answer choice C.), which is the decimal 0.2. 2.) You roll a number cube 20 times. The number 4 is rolled 8 times. What is the experimental probability of rolling a 4? A.) 40% <----- B.) 25% C.) 20% D.) 17%. (I multiplied 20x8=160, then divided by 4 which is 40 so I choose 40%.) 3.) The table below shows the results of flipping two coins. How does the experimental probability of getting at least one tails compare to the theoretical probability of getting at least one tails? HH=28, HT=22 TH=36, TT=16 A.) The experimental probability is 3% greater than the theoretical probability. <---- B.) The theoretical probability is 3% greater than the experimental probability. C.) The experimental probability is equal to the theoretical probability. D.) The experimental probability is 1% greater than the theoretical probability. (The theoretical probability is 25, and the experimental probability is 30. I choose answer A.)

satellite73 (satellite73):

hold the phone

satellite73 (satellite73):

the first one, you said you divided all that were not red by 50 that is the correct procedure what did you get?

OpenStudy (sugarbloom2004):

I got 0.6.

satellite73 (satellite73):

yeah that is right, but it is not C, C is 0.2

OpenStudy (sugarbloom2004):

Okay, so would it be my original answer response 0.6? I got 0.2 after someone told me to divide 0.6 by 3.

satellite73 (satellite73):

yes it is 0.6

OpenStudy (sugarbloom2004):

Okay thank you! Alright as for the second, and third question? Are my answer choices correct? I am not one-hundred percent sure on question number three.

satellite73 (satellite73):

You roll a number cube 20 times. The number 4 is rolled 8 times. What is the experimental probability of rolling a 4?

satellite73 (satellite73):

eight times out of twenty or \(\frac{8}{20}\) or reduce or turn in to a decimal

satellite73 (satellite73):

your explanation "I multiplied 20x8=160, then divided by 4 which is 40 so I choose 40%.) " not to be mean or anything, but it is not right first off, the number 40 is not 40% \(40\%=0.4=\frac{2}{5}\)

satellite73 (satellite73):

second, the number "4" in the question has nothing to do with it, it would be the same answer if they said, "roil a die 20 times the six come up 8"

satellite73 (satellite73):

ready for the next one?

OpenStudy (sugarbloom2004):

Yes.

satellite73 (satellite73):

HH=28, HT=22 TH=36, TT=16 how many rolls total?

OpenStudy (sugarbloom2004):

102 rolls.

satellite73 (satellite73):

and how many have "at least one tail"?

OpenStudy (sugarbloom2004):

Three.

satellite73 (satellite73):

oh on

satellite73 (satellite73):

\[HH=28, \color{red}{HT=22 ,TH=36, TT=16} \] all those have at least one tail

satellite73 (satellite73):

how many total have at least one tail?

OpenStudy (sugarbloom2004):

Three

satellite73 (satellite73):

maybe it is not clear what the numbers mean there

satellite73 (satellite73):

HH 28 means 28 times you got 2 heads, not one time

OpenStudy (sugarbloom2004):

Ohhh...

OpenStudy (sugarbloom2004):

Okay so I've got that one now. I understand the others, so now I just need help with three questions and then I am out of your hair. A coin is tossed and then a number cube is rolled. What is P(heads a number less than 5)? 1/3 5/12 2/3 5/6. <---- What is the probability of rolling a number cube and getting 4 and then an even number? 1/6 1/2 1/12 2/3<---- A bag of fruit contains 4 apples, 1 plum, 2 apricots, and 3 oranges. Pieces of fruit are drawn twice with replacement. What is P(apple, then apricot?) 4/5 2/25 3/25 3/5. <-----

satellite73 (satellite73):

A coin is tossed and then a number cube is rolled. What is P(heads a number less than 5)?

satellite73 (satellite73):

it is not \(\frac{5}{6}\)

satellite73 (satellite73):

how many numbers less than 5?

OpenStudy (sugarbloom2004):

There are three

satellite73 (satellite73):

name then

OpenStudy (sugarbloom2004):

Two, three, and four.

satellite73 (satellite73):

you missed one

OpenStudy (sugarbloom2004):

Oh and one.

satellite73 (satellite73):

so how many total?

OpenStudy (sugarbloom2004):

Four out of five or 4/5

satellite73 (satellite73):

how many sides does one die have?

OpenStudy (sugarbloom2004):

6.

satellite73 (satellite73):

right, so not four out of five try again

OpenStudy (sugarbloom2004):

Okay, so would it be two-thirds?

satellite73 (satellite73):

yes but that is not the final answer

OpenStudy (sugarbloom2004):

Okay so reducing it completely would be 1/3

satellite73 (satellite73):

what is the probability you toss heads on a coin

satellite73 (satellite73):

yes\[\frac{1}{2}\times \frac{2}{3}=\frac{1}{3}\]

satellite73 (satellite73):

What is the probability of rolling a number cube and getting 4 and then an even number?

satellite73 (satellite73):

what is the probability you roll a 4 on one die?

OpenStudy (sugarbloom2004):

1/6

satellite73 (satellite73):

yes, and the probability you roll an even number?

OpenStudy (sugarbloom2004):

3/6

satellite73 (satellite73):

yes multiply those two numbers \[\frac{1}{6}\times \frac{3}{6}\]

OpenStudy (sugarbloom2004):

It is 0.08% in a percentage and 1/12 as a fraction

satellite73 (satellite73):

yes

satellite73 (satellite73):

A bag of fruit contains 4 apples, 1 plum, 2 apricots, and 3 oranges. Pieces of fruit are drawn twice with replacement. What is P(apple, then apricot?)

satellite73 (satellite73):

find the probability you get an apple, then the probability you get a pear, then multiply

OpenStudy (sugarbloom2004):

There are no pears?

satellite73 (satellite73):

oops i means "apricot"

OpenStudy (sugarbloom2004):

Okay, the probability to get an apple would be 4/10, or 40% and the probability of getting an apricot would be 2/10 or 20%. With these, you would multiply 4/10x2/10, correct?

satellite73 (satellite73):

yes

OpenStudy (sugarbloom2004):

Alright, perfect! Thank you for your help. It is greatly appreciated. And how do you give a metal or do you automatically get that?

OpenStudy (sugarbloom2004):

Okay I figured it out, you have two medals.